A women has $34000 to invest. She invests some at 17% and the balance at 20%. Her total annual interest income is $6425. Fine the amount invested at each rate.

Answer by mananth(16949)

Part I 17.00% per annum ------------- Amount invested =x
Part II 20.00% per annum ------------ Amount invested = y =
34000
Interest----- 6425

Part I 17.00% per annum ---x
Part II 20.00% per annum ---y
Total investment
x + 1 y= 34000 -------------1
Interest on both investments
17.00% x + 20.00% y= 6425
Multiply by 100
17 x + 20 y= 642500.00 --------2
Multiply (1) by -17
we get
-17 x -17 y= -578000.00
Add this to (2)
0 x 3 y= 64500
divide by 3
y = 21500
Part I 17.00% $ 12500
Part II 20.00% $ 21500

CHECK
12500 --------- 17.00% ------- 2125.00
21500 ------------- 20.00% ------- 4300.00
Total -------------------- 6425.00

m.ananth@hotmail.ca